OpenStudy (anonymous):

Can someone help me with this proof.. If B is an orthonormal transform, y_1=Bx+2 and y_2=Bx_2 where x_i and y_i are vectors...

OpenStudy (anonymous):

If B is an orthonormal transform, y_1=Bx+2 and y_2=Bx_2 where x_i and y_i are vectors. Let \[|x|^{2} = x^{T} x = \sum_{ }^{ } x_{i} ^{2} \] show that \[|y_{1} - y_{2} |^{2} = |x_{1}- x_{2} |^{2} \]

OpenStudy (blockcolder):

I guess that's supposed to be \(y_1=Bx_1\)?

OpenStudy (anonymous):

Yeah sorry y_1=Bx_1

OpenStudy (blockcolder):

Was no vector space specified?

OpenStudy (anonymous):

Nope

OpenStudy (blockcolder):

I see. I'll rephrase \(Bx_1\) as \(B(x_1)\) so that it looks similar to function notation. Since B is an orthonormal transform, it is true that \[|x_1-x_2|^2=\left<x_1-x_2,x_1-x_2\right>=\left<B(x_1-x_2),B(x_1-x_2)\right>\]right?

OpenStudy (anonymous):

How did you get that to that conclusion?

OpenStudy (anonymous):

having a hard time grasping what orthonormal transform is

OpenStudy (blockcolder):

The definition of norm for arbitrary vector spaces is: \[|u|^2=\left<u,u\right>\] And an orthonormal transform \(B: V \to V\) is a transform such that \[\left<u,v\right>=\left<T(u),T(v)\right>\] Is this different from the definition that you are given?

OpenStudy (anonymous):

No no that's very similar to what im given

OpenStudy (anonymous):

Okay that make sense then

OpenStudy (anonymous):

but so you can insert a constant

OpenStudy (anonymous):

into the vector just like that?

OpenStudy (anonymous):

I don't get why it's B: V-> V. and then <u,v>=<T(u), T(v)>

OpenStudy (anonymous):

Oh you mean <B(u), B(v)>?

OpenStudy (blockcolder):

Yeah. Sorry bout that.

OpenStudy (anonymous):

Yeah no worries okay I get that step you wrote above now

OpenStudy (anonymous):

so what's next?

OpenStudy (blockcolder):

Anyway, we know that \(|x_1-x_2|^2=\left<B(x_1-x_2),B(x_1-x_2)\right>\). Use the fact that B is a linear transformation to rewrite the RHS as \(\left<B(x_1)-B(x_2),B(x_1)-B(x_2)\right>\). Finally, replace B(x_1) with y_1 and B(x_2) with y_2 then replace the inner product with the square of a norm.

OpenStudy (anonymous):

Oh I see got it hat make sense!

OpenStudy (anonymous):

but how do you know b is a linear transform

OpenStudy (blockcolder):

Because B is an orthonormal transform.

OpenStudy (anonymous):

lol okay I guess I have to review my terms..

OpenStudy (anonymous):

Thanks very much for your help!

OpenStudy (anonymous):

oh but so you don't utilize the transpose part they gave us?

OpenStudy (blockcolder):

Not really, since it wasn't specified that the vectors come from R^n.

OpenStudy (anonymous):

But if it was given do you know how'd you use it to prove it?

OpenStudy (anonymous):

the reason is this course does a lot of transposing..

OpenStudy (blockcolder):

If the specified vector space is R^n, and all of x_1, x_2, y_1, y_2 are column vectors, then you can assume that the standard matrix of B is an n-by-n orthogonal matrix. Call this matrix A. Since A is orthogonal, it is true that \(A^T=A^{-1}\). Thus, \(B(x_1)=Ax_1=y_1\) and \(B(x_2)=Ax_2=y_2\). This time, you start from \(|y_1-y_2|^2=(y_1-y_2)^T(y_1-y_2)=(y_1^T-y_2^T)(y_1-y_2)\) Then replace each y with Ax to get \(((Ax_1)^T-(Ax_2)^T)(Ax_1-Ax_2)\).

OpenStudy (blockcolder):

Simplify the first part of that last expression to get \[(x_1^TA^T-x_2^TA^T)=(x_1^T-x_2^T)A^T=(x_1^T-x_2^T)A^{-1}=(x_1-x_2)^TA^{-1}\] So that last one is simplified into \[(x_1-x_2)^TA^{-1}A(x_1-x_2)=(x_1-x_2)^TI(x_1-x_2)=(x_1-x_2)^T(x_1-x_2)\] And that last part should look familiar.

OpenStudy (anonymous):

Okay everything make sense except the last part, how were you able to just multply (x_1-x_2)^T A^-1 with A(x_1-x_2)?

OpenStudy (blockcolder):

This was the last expression 3 posts above: \[\color{red}{((Ax_1)^T-(Ax_2)^T)}(Ax_1-Ax_2)\] I then showed that \[\color{red}{((Ax_1)^T-(Ax_2)^T)=(x_1-x_2)^TA^{-1}}\] Thus, in the end, we get \[\color{red}{(x_1-x_2)^TA^{-1}}A(x_1-x_2)\]

OpenStudy (anonymous):

Ohhh I see what you did now!

OpenStudy (anonymous):

Thanks so much for your help, you have no idea how much this clarified things

OpenStudy (anonymous):

really appreciate it thanks a lot!

OpenStudy (anonymous):

sorry last question what property allows you do (y_1 - y_2) ^T to be (y_1^T - y_2 ^ T)?

OpenStudy (anonymous):

never mind I guess it's just a property it's in my notes. Okay thanks again

OpenStudy (blockcolder):

No problem. :)